A shipment of 11 printers contains 2 that are defective. Find the probability that a
sample of 4 printers, drawn from the 11, will not contain a defective printer.

Answer by stanbon(75887)

A shipment of 11 printers contains 2 that are defective. Find the probability that a sample of 4 printers, drawn from the 11, will not contain a defective printer.
----
Binomial Problem with n = 4 and p = 2/11
---
P(x = 0) = (9/11)^4 = 0.4481
============
Cheers,
Stan H.